A sample of caustic soda weighing 1.675 g was found to contain 97.25% of total alkali calculated as NaOH and of which is 2.48% was Na2CO3. What volume of 1.15 N acid could have been consumed in the Polypropylene and MO titrations as ordinarily performed?

MW: NaOH = 40.0, Na2CO3 = 106.0
1
Expert's answer
2020-10-20T08:35:25-0400

(Assuming that 1.15 N actually means 1.15 M HCl acid, Polypropylene was meant to be Phenolphthalein indicator and MO is Methyl Orange indicator.)

Mass of NaOH "=\dfrac{97.25}{100}x1.475 = 1.4344375g"


Mass of Na2CO3"\dfrac{2.48}{100}x1.475 = 0.03658g"


Moles of NaOH "=\dfrac{1.4344375}{40} = 0.03586mol"


Moles of Na2CO3"=\dfrac{0.03658}{106} = 3.383x10^-4mol"


Moles of acid required to react with all the NaOH

Mole ratio=1:1

Hence moles of acid "=\dfrac{1}{1}x0.03586 = 0.03586 mol"


Moles of acid required to react with all Na2CO3

Mole ratio = 2:1

Hence moles of acid"=\dfrac{1}{2}x3.383x10^-4 = 1.692x10^-4 mol"


Total moles of acid used = 0.03586 + 1.692x10^-4 = 0.03602915mol.


Volume of acid"=\dfrac{moles x 1000}{molarity} = \dfrac{0.0360291x1000}{1.15}"


= 31.33 cm3 for Phenolphthalein


For MO"=\dfrac{31.33}{100} = 0.3133" cm3